Calculate the energy required in joules to raise the temperature of 450 grams of water from 15°C to 85°C. (The specific heat capacity of water is 4.18 J/g/°C)
1
Expert's answer
2019-09-27T09:43:44-0400

The energy required in joules to raise the temperature of 450 grams of water from 15°C to 85°C:


"Q=mc(T_2-T_1)"

"Q=(450)(4.18)(85-15)=132000\ J"